OpenStudy (anonymous):

please tell me if im right on this one f(x)= x^2+2x+24 find the x and y intercepts x=(-4,6) y= 0,10)

OpenStudy (anonymous):

x intercept has to have the y coordinate 0 and the y intercept has to have x coordinate 0

OpenStudy (anonymous):

to find x intercept, you'll set y=x^2+2x+24=0 Since this equation x^2+2x+24 is always positive, for any x, there is no x intercept The graph of equation is above x axis

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

so there is no x intercept?

OpenStudy (anonymous):

it is simple, the graph of the function y=x^2+2x+24 is a parabola that is located above x axis and it does not intercept x axis ever

OpenStudy (anonymous):

Definetely, there is no x intercepts, what so ever

OpenStudy (anonymous):

parabola is intercepting only y axis at the point (0,24)